OpenStudy (anonymous):

Subtract 5 2/5 -2 9/10. Justify your steps.

OpenStudy (zzr0ck3r):

\[5 \frac{2}{5}-2 \frac{9}{10}?\]

OpenStudy (anonymous):

i know the answer i just dont know the steps to get to it

OpenStudy (zzr0ck3r):

im asking if I read you question right

OpenStudy (anonymous):

yes you did

OpenStudy (zzr0ck3r):

ok we want to get the fraction to look alike so multiply the first fraction by 2 on top and bottom \[5\frac{2}{5}-2\frac{9}{10}=5\frac{2}{5}*\frac{2}{2}-2\frac{9}{10}=5\frac{4}{10}-2\frac{9}{10}\] we want to have positive outcome when we divide the fractions we borrow 1 from the 5 \[5\frac{4}{10}= 4\frac{14}{10}\] so we have \[4\frac{14}{10}-2\frac{9}{10}=2\frac{5}{10}\]

OpenStudy (zzr0ck3r):

that should say we want positive outcome when we subtract the fractions

OpenStudy (texaschic101):

are you understanding what zzr0ck3r is saying ?

OpenStudy (anonymous):

Yes i do and thank you so much

OpenStudy (texaschic101):

no problem...I was just checking :)